A machine’s motion is as shown below. NOTE: The line is NOT straight between t=6 and t=8s.

1) Draw the s-t graph and a-t graph. Label key points, axes; clearly distinguish between straight and curved lines. Enter the value of acceleration, a, at t=8 s, for the online answer below.

2) Determine the function of acceleration (a) and distance (s) in terms of time (t) over the time interval 6 to 8 seconds.

18 15 12 9 6 velocity (ft/s) v = 36-0.52 (from t=6 to t=8) 0 -6 -9 a time (5)
Answer #1

a Реле, For oct 22 velocity varies linearly, V-oto -9 2 ds 4.5t 2 Hica (-4,54 - = - 4.5 o 4.5 (23² d=2 t=2 For 2 ef 55 » velo(4.5 (5² – 27 (5) -[(4,5) 12² - 27 (2) (-9) = - 22.5+ 36 = 13.5 = 4.5 m ft 4.5 m ft For 52+26 velocity r= constant av = 18 .3. s- & graph in + For oct ca Subs - 4.5$ Z 45.13 45 До 35 & For 25t5 30 27.5 SU) = 4,51 - 27t-g 20 15 10 (At) 4.5 For 54fc64 Q1.2 Function of accelestion over 6ct as Ha=rt k Angoes 7 361 – 6, EP 72,5 Borse 3